OpenStudy (anonymous):

Which point lies on the graph of the following piecewise function? y = x – 1 if x < 2 y = x^2 if x greater than or equal to 2 a.(1, 1) b.(2, 1) c.(–1, 1) d.(3, 9)

OpenStudy (anonymous):

check and see. for example, if x = 1 since 1< 2 you would use the first rule and take 1-1 = 0, so (1,0) is on the graph, not (1,1)

OpenStudy (anonymous):

if x = 2 you would use the second rule, so you would get y = 2^2 = 4 and therefore (2,4) is on the graph, not (2,1)

OpenStudy (anonymous):

i think you can skip down to (3,9) and see that since 3 > 2 you use the second rule and put y = 3^2 = 9, so (3, 9) is on the graph

OpenStudy (anonymous):

Thank you, so you wouldn't use 2 for the equation f(x) = x - 1 if x < 2?

OpenStudy (anonymous):

no since 2 is not less than 2 you would use y = 2^2